A 645-N man rides an elevator. If 250.0 W is needed to lift a person 6.00 m, how long will it take for the elevator to lift the man?


1
Expert's answer
2021-02-22T10:20:01-0500

By the definition of the power, we have:


"P=\dfrac{W}{t}=\dfrac{mgh}{t},""t=\dfrac{mgh}{P}=\dfrac{645\ N\cdot6.0\ m}{250\ W}=15.5\ s."